# Hierarchical models for small area estimation using zero-inflated forest inventory variables: comparison and implementation

### v1

@@ Abstract
National Forest Inventory (NFI) data are typically limited to sparse networks of sample locations due to cost constraints. While traditional design-based estimators provide reliable forest parameter estimates for large areas, there is increasing interest in model-based small area estimation (SAE) methods to improve precision for smaller spatial, temporal, or biophysical domains. SAE methods can be broadly categorized into area- and unit-level models, with unit-level models offering greater flexibility—making them the focus of this study. Ensuring valid inference requires satisfying model distributional assumptions, which is particularly challenging for NFI variables that exhibit positive support and zero inflation, such as forest biomass, carbon, and volume. Here, we evaluate a class of two-stage unit-level hierarchical Bayesian models for estimating forest biomass at the county-level in Washington and Nevada, United States. We compare these models to simpler Bayesian single-stage and two-stage frequentist approaches. To assess estimator performance, we employ simulated populations and cross-validation techniques. Results indicate that small area estimators that incorporate a two-stage approach to account for zero inflation, county-specific random intercepts and residual variances, and spatial random effects provide the most reliable county-level estimates. Additionally, findings suggest that unit-level cross-validation within the training dataset is as effective as area-level validation using simulated populations for model selection. We also illustrate the usefulness of simulated populations for better assessing qualities of the various estimators considered.
Keywords: model-based, Gaussian process, spatial, Bayesian, carbon, biomass, simulation
@@ 1 Introduction
National Forest Inventories (NFIs) play a critical role in collecting data and monitoring forest trends to assess resource availability, health, composition, and other economic and ecological attributes across various spatial scales within a given country. In the United States, the NFI is conducted by the United States Department of Agriculture (USDA) Forest Service through the Forest Inventory and Analysis (FIA) Program. Traditionally, NFIs such as FIA have been designed to provide precise estimates at broader spatial scales, such as state-level assessments of forest attributes like timber volume and biomass. However, there is growing national interest, along with increased funding, in obtaining more precise biomass estimates at finer spatial scales, such as the county-level (Wiener et al.,, 2021; Prisley et al.,, 2021; U.S. Senate,, 2023). This rising demand, coupled with the widespread availability of high-resolution remote sensing data, has prompted researchers to develop and apply innovative small area estimation (SAE) methods that integrate FIA data with remote sensing products (Cao et al.,, 2022; May et al.,, 2023; Finley et al.,, 2024).
@@
The advantages and trade-offs of SAE approaches have only begun to be explored in the forest inventory literature. Area-level modeling often benefits from a more linear relationship between response and predictor variables and does not require precise plot locations, which is particularly useful given the often confidentiality of NFI plot data. However, aggregation leads to data loss, limits the ability to model fine-scale (i.e., unit-level) relationships, precludes delineation of new areas of interest after model fitting, and imposes statistical assumptions that might be difficult to justify. For example, in the classical Fay-Herriot model, within-area variance of the direct estimate is assumed to be fixed and known, although in practice it is estimated from limited data. These variances enter the area-level model through a random effect, often without strong theoretical justification or consistency between the two inferential paradigms. In contrast, unit-level approaches leverage precise plot locations to model fine-scale spatial relationships more effectively, making them particularly valuable when such data are available.
Unlike design-based estimators, which are determined by the sampling design, inference from model-based estimators relies entirely on the selection of an appropriate model. Consequently, we must take special care when specifying SAE models and conduct rigorous model checking. One of the most effective ways to assess SAE models is through simulated populations that closely resemble the true, but only partially observed, population of interest. Simulated populations allow us to explore how inference varies under different conditions (e.g., varying sample sizes) and to compare our estimates against “true” values. To ensure a meaningful evaluation, we must generate these simulated populations using methods that are not similar to the models we hope to assess.
Beyond assessment using simulated populations, we can evaluate models through cross-validation using observed data (e.g., leave-one-out or -fold cross-validation). However, in SAE studies, the primary parameters of interest exist at the area level. Unit-level models can be assessed using cross-validation at the unit level (i.e., iteratively holdout one or more observations, predict for those holdout, and compare the predictions to the holdout true values); however, such unit-level assessments might not reveal how well the estimator performs once predictions are aggregated to the desired areas of interest.
@@
In forest inventory applications, zero-inflation has received relatively limited attention. Finley et al., (2011) developed a two-stage model for zero-inflation in continuous forest attributes such as biomass, volume, and age, employing a hierarchical Bayesian framework with Gaussian process-based spatial random effects. Their approach enables unit-level predictions of forest attributes along with uncertainty quantification, though they did not directly produce small area estimates. More recently, White et al., (2025) applied the zero-inflated SAE model from Chandra and Sud, (2012) to FIA data in Nevada, generating county-level biomass estimates. Their study compared the zero-inflated estimator to other commonly used small area estimators, including the Battese-Harter-Fuller unit-level model and the Fay-Herriot area-level model (Fay and Herriot,, 1979; Battese et al.,, 1988). Their simulation results indicate the zero-inflated model improves point estimates and produces competitive MSE estimates, though further refinements remain possible (see Figure 2 in White et al.,, 2025).
@@
To evaluate these nine estimators, we conduct a simulation study based on the methodology of White et al., 2024a , detailed in Section 2.6. Section 3 presents the results of the simulation study and applies the estimators to FIA data, with cross-validation used to assess model performance at the unit level. Finally, we summarize our findings, discuss their implications, and suggest directions for future research in Section 4.
@@ 2.1 Data
The motivating data are from the USDA Forest Service FIA program and comprise inventory plot measurements of live aboveground tree biomass density (Mg/ha). These data were drawn from the most current sampled measurement for each plot in the FIA database downloaded on February 8, 2023 for the states of Washington and Nevada (Burrill et al.,, 2023). Washington was selected because it has large differences in biomass across counties, ranging from massive biomass densities on the Olympic Peninsula to near zero biomass in counties east of the Cascade Mountain Range. Nevada was selected because it has a fairly unique distribution of forest biomass, where much of the state’s arid environment has little to no biomass which is punctuated with sky islands where there is non-zero forest biomass. Each state has approximately ten years of FIA data, ending in year 2019, that were derived from a panel of plots measured annually across a systematic sample of hexagons approximately 2,500 hectare in size. Biomass values were from live trees only and include all trees 1.0 inch diameter and greater.
@@
Estimators and simulations, described in Sections 2.2 and 2.6, respectively, were informed using five auxiliary variables: National Land Cover Dataset Analytical Tree Canopy Cover 2016 (hereafter tcc); LANDFIRE 2010 Digital Elevation Model (hereafter elev); US Geological Survey Terrain Ruggedness Index (hereafter tri); PRISM mean annual precipitation, 30yr normals (1991-2020) (hereafter ppt); and LANDFIRE 2014 tree/non-tree lifeform mask (hereafter tnt) (Yang et al.,, 2018; U.S. Geological Survey,, 2019; Daly et al.,, 2002; Rollins,, 2009; Picotte et al.,, 2019). The tcc variable is a measure of average tree canopy cover in a given pixel, the elev variable gives the elevation at a given pixel, the tri variable gives the terrain ruggedness at a given pixel, the ppt variable is a measure of average precipitation at a given pixel over 30 years, and the tnt variable is a binary variable distinguishing between pixels with and without trees. These auxiliary variables were resampled to 90 meter resolution and available wall-to-wall in both states. At locations with FIA plots these variables are matched with the corresponding plot and then used as predictors in the estimators’ regression component and to inform simulated population generation. The use of these variables in different model types and simulated population generation is shown in Table 1.
@@ 2.2 Model-based estimation
Nine candidate model-based estimators are used to estimate the average biomass at the county level across Nevada and Washington. The first estimator follows a frequentist mode of inference and is constructed using a two-stage regression. The eight additional estimators use a Bayesian mode of inference and are constructed using one- and two-stage regressions. A brief description of the candidate estimators is provided in Table 2.
@@
| B ZI CVC CRV | The same as B ZI CVC, but with county-specific residual variance. |
@@
Models are fit at the unit level, which sets plot-level biomass (Mg/ha) as the response variable. To better meet subsequent models’ assumption of normally distributed residuals and to ensure positive support for predictions, models are fit to a transformed response variable. Often in such settings the logarithm is a natural choice; however, we found the logarithm to be too strong, resulting in a skewed response distribution once zeros were accommodated. We prefer root function transformations that were chosen to be state specific, with stronger roots for high biomass states and weaker roots for low to moderate biomass states. Specifically, we used a fourth and square root transformation for Washington and Nevada, respectively.
Despite our models being fit at the unit level, our inferential goal is to estimate average biomass per hectare at the county level, which we denote where indexes county within a given state. In subsequent sections we present the candidate unit-level models and how each is used to estimate the small area parameters of interest.
2.3 Frequentist two-stage estimator
Here we develop a frequentist two-stage approach to estimation (F ZI CVI). This estimator and its MSE estimator were introduced by Chandra and Sud, (2012) and later applied to forest inventory data by White et al., (2025) with promising results in the state of Nevada. At generic spatial location the transformed non-zero forest biomass is modeled as
@@ 2.3 Frequentist two-stage estimator
2.4 Bayesian estimators
Here we consider a class of hierarchical Bayesian models. These models are primarily two-stage estimators similar to the frequentest two-stage estimator developed in Section 2.3. For comparison, we also include some simpler, single-stage, models that do not explicitly accommodate excess zero values in the response.
@@ 2.4 Bayesian estimators
The first level of hierarchy is the model for . In our case, we only consider two options for estimation of : first, setting to 1; and second, using a linear mixed model with logit link function. In the second case, for simplicity, we only use one model for estimation of across all Bayesian estimators. Here, a realization of indicates whether or not the location is predicted to have biomass (1) or not (0). Then, we pass the realization of into the second level of hierarchy where it determines the expression of the mean and the associated variance term. In the subsequent development, is a given linear mixed model with county-varying intercept (CVI), county-varying coefficients (CVC), or space-varying intercept (SVI) components (Table 2). We implement and compare four different model forms for . The residual variance is estimated through a parameter when is 1, and set to a small value via a constant, , when is 0 (see Finley et al.,, 2011, Section 3 for details). For some candidate models we allow for county-specific residual variance (CRV) terms via county-specific s (Table 2).
@@
The first stage model used for all two-stage models is a generalized linear mixed model with Bernoulli response and a county-varying intercept defined as
where denotes the probability of non-zero response, and parameter and hyperparameter distributions defined as follows, , is the county specific random effect, i.e., when is in the th county, , and . All hyperparameters were set to induce non-informative prior distributions.
@@
where is a spatial random effect that adjusts the intercept based on residual spatial dependence. Here we estimate using a GP approximation called the Nearest Neighbor Gaussian Process (NNGP; Datta et al., 2016; Finley et al., 2019) that provides substantial improvements in run time, with negligible differences in inference and prediction, compared to a model that uses a full Gaussian Process. In brief, for this specification, the vector of random effects collected over locations is distributed multivariate normal with mean zero and covariance matrix that captures the spatial dependence among random effects, i.e., , where is the spatial variance and is the NNGP-derived correlation matrix that depends on a spatial correlation function, which in our case is an exponential, and decay parameter used to estimate the strength of correlation between any two locations. As with other variance parameters, we assume with hyperparameters set to induce a non-informative prior distribution. The spatial decay parameter is assumed to follow a uniform distribution with broad, non-informative, spatial support.
@@ 2.5.1 Frequentist two-stage estimator
Model parameters for two-stage frequentist model (F ZI CVI) described in Section 2.3 were estimated using restricted maximum likelihood via the R (R Core Team,, 2024) saeczi package (Yamamoto et al.,, 2025) which implements methods presented by Chandra and Sud, (2012).
@@
where the hat indicates each parameter’s maximum likelihood point estimate. The estimate for is then the average product of these predictions over the grid of prediction locations
where is the inverse of the transformation function used when fitting the model, allow us to revert back to biomass scale, and is the set of prediction locations within th county. This estimator’s MSE estimator comes from a parametric bootstrap introduced in Chandra and Sud, (2012) and discussed and explored in White et al., (2025).
@@ 2.5.2 Bayesian estimators
Parameter inference for Bayesian models described in Section 2.4 was based on Markov chain Monte Carlo (MCMC) samples from posterior distributions. Gibbs and Metropolis Hastings algorithms were implemented in C++ to efficiently samples from parameter posterior distributions. Code, additional information about the sampling algorithms, and example analyses using simulated data are given in Finley, (2025) and a list of prior distributions and hyperparameter values is given in Appendix A. Posterior inference is based on =3,000 post-convergence and thinned samples from three MCMC chains, i.e., 1,000 from each chain. We use convergence diagnostics and thinning rules outlined in Gelman et al., (2013).

### v2

@@ Abstract
National Forest Inventory (NFI) data are typically limited to sparse networks of sample locations due to cost constraints. While traditional design-based estimators provide reliable forest parameter estimates for large areas, there is increasing interest in model-based small area estimation (SAE) methods to improve precision for smaller spatial, temporal, or biophysical domains. SAE methods can be broadly categorized into area- and unit-level models, with unit-level models offering greater flexibility—making them the focus of this study. Ensuring valid inference requires satisfying model distributional assumptions, which is particularly challenging for NFI variables that exhibit positive support and zero inflation, such as forest biomass, carbon, and volume. Here, we evaluate a class of two-stage unit-level hierarchical Bayesian models for estimating forest biomass at the county-level in Washington and Nevada, United States. We compare these models to simpler Bayesian single-stage and two-stage frequentist approaches. To assess estimator performance, we employ simulated populations and cross-validation techniques. Results indicate that small area estimators that incorporate a two-stage approach to account for zero inflation, county-specific random intercepts and residual variances, and spatial random effects provide the most reliable county-level estimates. We illustrate the usefulness of simulated populations and cross-validation for assessing qualities of the various estimators considered.
Keywords: model-based estimation, Gaussian processes, Bayesian, forest biomass, simulation study
@@ 1 Introduction
National Forest Inventories (NFIs) play a critical role in collecting data and monitoring forest trends to assess resource availability, health, composition, and other economic and ecological attributes across various spatial scales within a given country. In the United States, the NFI is conducted by the United States Department of Agriculture (USDA) Forest Service through the Forest Inventory and Analysis (FIA) Program. Traditionally, NFIs such as FIA have been designed to provide precise estimates at broader spatial scales, such as state-level assessments of forest attributes like timber volume and biomass. However, there is growing interest among data users, along with increased funding, in obtaining more precise biomass estimates at finer spatial scales, such as the county-level (Prisley et al.,, 2021; Wiener et al.,, 2021; U.S. Senate,, 2023). This rising demand, coupled with the widespread availability of high-resolution remote sensing data, has prompted researchers to develop and apply innovative small area estimation (SAE) methods that integrate FIA and similar NFI data with remote sensing products (Cao et al.,, 2022; May et al.,, 2023; Finley et al.,, 2024).
@@
The advantages and trade-offs of SAE approaches have only begun to be explored in the forest inventory literature. Area-level modeling often benefits from a more linear relationship between response and predictor variables and does not require precise plot locations, which is particularly useful given that NFI plot locations are often confidential. However, aggregation leads to data loss, limits the ability to model fine-scale (i.e., unit-level) relationships, precludes delineation of new areas of interest after model fitting, and imposes statistical assumptions that might be difficult to justify. For example, in the classical area-level Fay-Herriot model, within-area variance of the direct estimate is assumed to be fixed and known, although in practice it is estimated from limited data (Fay and Herriot,, 1979; Rao and Molina,, 2015). These variances enter the area-level model through a random effect, often without strong theoretical justification or consistency between the two inferential paradigms (see, e.g., Shannon et al.,, 2024, for discussion). In contrast, unit-level approaches leverage precise plot locations to model fine-scale spatial relationships more effectively, making them particularly valuable when such data are available.
Unlike design-based estimators, which derive their properties from the sampling design, model-based estimators rely entirely on the assumed data-generating process and the selection of an appropriate model. As a result, particular care must be taken in specifying SAE models and conducting rigorous model checking. One of the most effective approaches to evaluating SAE models is through the use of simulated populations that closely mimic the true—but only partially observed—population of interest. Simulated populations enable us to examine how inference varies under different conditions (e.g., varying sample sizes) and to compare our estimates against “true” values. To ensure that this evaluation is meaningful, it is essential to generate simulated populations using processes that are not closely aligned with the models under assessment.
Beyond assessment using simulated populations, we can evaluate models through cross-validation using observed data (e.g., leave-one-out or -fold cross-validation). However, in SAE studies, the primary parameters of interest exist at the area level. Unit-level models can be assessed using cross-validation at the unit level (i.e., iteratively holdout one or more observations, predict for those holdout, and compare the predictions to the holdout true values); however, we must be careful to verify the unit-level assessments align with how well the estimator performs once predictions are aggregated to the desired areas of interest.
@@
In forest inventory applications, zero-inflation has received relatively limited attention. Finley et al., (2011) developed a two-stage model for zero-inflation in continuous forest attributes such as biomass, volume, and age, employing a hierarchical Bayesian framework with Gaussian process-based spatial random effects. Their approach enables unit-level predictions of forest attributes along with uncertainty quantification, though they did not directly produce small area estimates. More recently, White et al., (2025) applied the zero-inflated SAE model from Chandra and Sud, (2012) to FIA data in Nevada, generating county-level biomass estimates. Their study compared the zero-inflated estimator to other commonly used small area estimators, including estimators based on the Battese-Harter-Fuller unit-level model and the Fay-Herriot area-level model (Fay and Herriot,, 1979; Battese et al.,, 1988). Their simulation results indicate the zero-inflated estimator improves point estimates and produces competitive MSE estimates, though further refinements remain possible (see Figure 2 in White et al.,, 2025).
@@
To evaluate the nine estimators introduced in Section 2, we conduct a simulation study following the methodology of White et al., 2024a , as described in Section 2.7. Section 3 presents the simulation results and applies the estimators to FIA data, using cross-validation to assess unit-level model performance. Finally, Section 4 summarizes our findings, discusses their implications, and outlines directions for future research.
@@ 2.1 Data
The motivating data are from the USDA Forest Service FIA program and comprise inventory plot measurements of live aboveground tree biomass density (Mg/ha). These data were drawn from the most current sampled measurement for each plot in the FIA database downloaded on February 8, 2023 for the states of Washington and Nevada (Burrill et al.,, 2023). Washington was selected because it has large differences in biomass across counties, ranging from massive biomass densities on the Olympic Peninsula to near zero biomass in counties east of the Cascade Mountain Range. Nevada was selected because it has a fairly unique distribution of forest biomass, where much of the state’s arid environment has little to no biomass which is punctuated with sky islands where there is non-zero forest biomass. Each state has approximately ten years of FIA data, ending in year 2019, that were derived from a panel of plots measured annually across a systematic sample of hexagons approximately 2,500 hectares in size. Biomass values were from live trees only and include all trees 1.0 inch diameter and greater.
@@
Estimators and simulations, described in Sections 2.3 and 2.7, respectively, were informed using five auxiliary variables: National Land Cover Dataset Analytical Tree Canopy Cover 2016 (hereafter tcc); LANDFIRE 2010 Digital Elevation Model (hereafter elev); US Geological Survey Terrain Ruggedness Index (hereafter tri); PRISM mean annual precipitation, 30yr normals (1991-2020) (hereafter ppt); and LANDFIRE 2014 tree/non-tree lifeform mask (hereafter tnt) (Daly et al.,, 2002; Rollins,, 2009; Yang et al.,, 2018; Picotte et al.,, 2019; U.S. Geological Survey,, 2019). The tcc variable is a measure of average tree canopy cover in a given pixel, the elev variable gives the elevation at a given pixel, the tri variable gives the terrain ruggedness at a given pixel, the ppt variable is a measure of average precipitation at a given pixel over 30 years, and the tnt variable is a binary variable distinguishing between pixels with and without trees. These auxiliary variables were resampled to 90 meter resolution and available wall-to-wall in both states. At locations with FIA plots these variables are matched with the corresponding plot and then used as predictors in the models’ regression components and to inform simulated population generation.
2.2 Variable Selection
The five auxiliary variables described in Section 2.1 were used at multiple stages of analysis, including the generation of simulated populations for Washington and Nevada, the fitting of model-based estimators on these simulated populations, and the application of these estimators to FIA data.
For simulated population generation, variable selection was informed by domain knowledge about regional biomass drivers. In both Washington and Nevada, tcc and elev were selected as core predictors due to their expected influence on biomass distribution. In Nevada, we additionally included tri, given the ecological relevance of sky islands in structuring biomass presence. In Washington, ppt was added to reflect the east–west precipitation gradient across the Cascade Range, which significantly influences biomass patterns. To ensure distinct imputation strategies for different vegetation conditions, we stratified the simulated populations in both states by tnt.
For the model-based estimators applied to both simulated and FIA data, we used the same predictor variables within each state to maintain consistency with the corresponding data-generating processes. Initially, the tnt variable was excluded from model-based estimators. While this exclusion had little effect in Nevada, omitting tnt in Washington led to poor predictive performance in the Bernoulli models. As a result, we included tnt as a predictor in all Bernoulli models for Washington.
Table 1 summarizes the use of auxiliary variables across model types and data sources.
@@ 2.3 Model-based estimation
We evaluate nine candidate model-based estimators for estimating average biomass density at the county level across Nevada and Washington. The first estimator employs a frequentist approach using a two-stage regression. The remaining eight adopt a Bayesian framework, incorporating both one-stage and two-stage regression structures. A summary of these estimators is provided in Table 2.
All models are fit at the unit level, with plot-level biomass serving as the response variable. To improve adherence to normality assumptions and to ensure positive support for predicted values, we apply a transformation to the response. While logarithmic transformations are common in such contexts, we found them overly aggressive in this application, especially when accommodating zero biomass values, which resulted in highly skewed transformed distributions. Instead, we apply root-based transformations, selecting state-specific powers to match regional biomass distributions: a fourth-root transformation for Washington and a square-root transformation for Nevada.
Although the models are fit at the unit level, our primary inferential target is the average biomass per hectare at the county level, denoted , where indexes counties within each state. In the sections that follow, we describe the candidate unit-level models in detail and outline how each is used to estimate these small area parameters of interest.
@@
| B ZI CVC CRV | The same as B ZI CVC, but with county-specific residual variances. |
@@
2.4 Models for frequentist two-stage estimation
Here we introduce the models used to develop a frequentist two-stage approach to estimation (F ZI CVI). The resultant estimator and its MSE estimator were introduced by Chandra and Sud, (2012) and later applied to forest inventory data by White et al., (2025) with promising results in the state of Nevada. At generic spatial location the transformed non-zero forest biomass is modeled as
@@ 2.4 Models for frequentist two-stage estimation
2.5 Models for Bayesian estimation
Here we consider a class of hierarchical Bayesian models. These models are primarily two-stage models similar to the frequentest two-stage model developed in Section 2.4. For comparison, we also include some simpler, single-stage, models that do not explicitly accommodate excess zero values in the response.
@@ 2.5 Models for Bayesian estimation
The first level of hierarchy is the model for . In our case, we only consider two options for estimation of : first, setting to 1; and second, using a linear mixed model with logit link function. In the second case, for simplicity, we only use one model for estimation of across all Bayesian estimators. Here, a realization of indicates whether or not the location is predicted to have biomass (1) or not (0). Then, we pass the realization of into the second level of hierarchy where it determines the expression of the mean and the associated variance term. In the subsequent development of models, is a given linear mixed model with county-varying intercept (CVI), county-varying coefficient (CVC), or space-varying intercept (SVI) components (Table 2), resulting in four different model forms for that we implement and compare. The residual variance is estimated through a parameter when is 1, and set to a small value via a constant, , when is 0 (see Finley et al.,, 2011, Section 3 for details). For some candidate models we allow for county-specific residual variance (CRV) terms via county-specific s (Table 2).
@@
[…]
where denotes the probability of non-zero response, and parameter and hyperparameter distributions are defined as follows, , is the county specific random effect, i.e., and when is in the th county, , and . All hyperparameters were set to induce non-informative prior distributions.
@@
where is a spatial random effect that adjusts the intercept based on residual spatial dependence. Here we estimate using a Gaussian Process approximation called the Nearest Neighbor Gaussian Process (NNGP; Datta et al., 2016; Finley et al., 2019) that provides substantial improvements in run time, with negligible differences in inference and prediction, compared to a model that uses a full Gaussian Process. In brief, for this specification, the vector of random effects collected over locations is distributed multivariate normal with mean zero and covariance matrix that captures the spatial dependence among random effects, i.e., , where is the spatial variance and is the NNGP-derived correlation matrix that depends on a spatial correlation function, which in our case is exponential, and decay parameter used to estimate the strength of correlation between any two locations. As with other variance parameters, we assume with hyperparameters set to induce a non-informative prior distribution. The spatial decay parameter is assumed to follow a uniform distribution with broad, non-informative, spatial support.
@@ 2.6.1 Frequentist two-stage estimator
Model parameters for the two-stage frequentist model (F ZI CVI) described in Section 2.4 were estimated using restricted maximum likelihood via the R (R Core Team,, 2024) saeczi package (Yamamoto et al.,, 2025) which implements methods presented by Chandra and Sud, (2012).
@@
where the indicates each parameter’s maximum likelihood point estimate. The estimate for is then the average product of these predictions over the grid of prediction locations
where is the inverse of the transformation function used when fitting the model, allowing us to revert back to biomass scale, and is the set of prediction locations within the th county. This estimator’s MSE estimator comes from a parametric bootstrap introduced in Chandra and Sud, (2012) and discussed and explored in White et al., (2025).
@@ 2.6.2 Bayesian estimators
Parameter inference for Bayesian models described in Section 2.5 was based on MCMC samples from posterior distributions. Gibbs and Metropolis Hastings algorithms were implemented in C++ to efficiently sample from parameter posterior distributions. Code, additional information about the sampling algorithms, and example analyses using simulated data are given in Finley, (2025) and a list of prior distributions and hyperparameter values is given in Appendix A. Posterior inference is based on =3,000 post-convergence and thinned samples from three MCMC chains, i.e., 1,000 from each chain. We use convergence diagnostics and thinning rules outlined in Gelman et al., (2013).